A machine uses 2.5 litres of fuel each hour it runs. Its fuel tank was filled with 10 L, but 1.5 L have already been used. How many more hours will the machine run?
1
Expert's answer
2019-10-28T11:40:24-0400

"t=\frac{10-1.5}{2.5}=\frac{8.5}{2.5}=3.4" hours.